OpenStudy (anonymous):

find the tangent of the ellipse \(\large \frac{x^2}{4}+y^2=1\) at the point \(p (\sqrt 2, \frac{1}{\sqrt2})\)

OpenStudy (anonymous):

@ganeshie8 @phi @jdoe0001 @myininaya

myininaya (myininaya):

did you find y'?

OpenStudy (anonymous):

\(\large y=\sqrt (1-\frac{x^2}{x}) \)

myininaya (myininaya):

why didn't you use implicit differentiation?

OpenStudy (anonymous):

idk, u told me "did you find y'? " so i thought of finding y :O

myininaya (myininaya):

do you know how to do implicit differentiation? you can solve for y and find y' explicitly.. it just looks uglier.

OpenStudy (anonymous):

could u plz show it ?

OpenStudy (jdoe0001):

am I wrong by thinking -> \(\bf \cfrac{x^2}{4}+y^2=1\implies x^2+4y^2-4=0\quad ?\)

OpenStudy (jdoe0001):

hmmm I see... nevermind that =)

myininaya (myininaya):

Here is example to find y': \[x^2+xy^2=1+yx \] Take derivative of both sides with respect to x \[\frac{d}{dx}(x^2+xy^2)=\frac{d}{dx}(1+yx)\] Use the sum rule for differentiation \[\frac{d}{dx}(x^2)+\frac{d}{dx}(xy^2)=\frac{d}{dx}(1)+\frac{d}{dx}(yx)\] We know how to find derivative of x^2, it is 2x now for finding derivative of (xy^2) or (yx) we need to know the product rule since we have a product \[2x+y^2 \frac{d}{dx}(x)+x \frac{d}{dx}(y^2)=\frac{d}{dx}(1)+y \frac{d}{dx}(x)+x \frac{d}{dx}y\] now I'm going to go ahead and replace the derivative of 1 with 0 since we know derivative of any constant is 0 now for find derivative of y^2 or y we need to know chain rule (y^2)'=2y*y' <---remember y=y(x) (y)'=y' so we have now \[2x+y^2+2yy'=y+xy'\] Then you can solve this pretty easily for y. By putting your terms that have a y' together and put everything that doesn't have a y' on the other side of the equation like this: \[2yy'-xy'=y-2x-y^2 \] We do this so we can factor out the y' and divide both sides by what y' is being multiplied by to obtain what y' is like so: \[y'(2y-x)=y-2x-y^2 => y'=\frac{y-2x-y^2}{2y-x}\]

OpenStudy (anonymous):

btw , i was thought of r(t)=[2 cos t , sin t ] i sow it in the text book but dnt know what it is :\ then it show r'(t)=[- sin t, cos t)

OpenStudy (anonymous):

thax @myininaya its awesome !!

myininaya (myininaya):

ok you try differentiating x^2/4+y^2=1 to find y' now

myininaya (myininaya):

\[\frac{d}{dx}(\frac{x^2}{4})=?\]

OpenStudy (anonymous):

1/2 x ?

myininaya (myininaya):

ok good now \[\frac{d}{dx}(y^2)=?\]

myininaya (myininaya):

I actually already told you this one..

OpenStudy (anonymous):

yeah ! im still reading ur sol ! its a bit long , but thx for detail :) i cant understand it nw thank you :D

myininaya (myininaya):

ok well do you know how to find derivative of (4x-1)^2?

OpenStudy (anonymous):

8(4x-1)

myininaya (myininaya):

and you did that by finding derivative of the inside times the the derivative of the outside so like you did this: 2(4x-1)*(4)

myininaya (myininaya):

well what if y=4x-1 and y'=4 so we have derivative of y^2 is 2(y)*y' right?

OpenStudy (anonymous):

right

myininaya (myininaya):

so when we differentiate x^2/4+y^2=1 what do we get as the very next step

OpenStudy (anonymous):

y2=1- x^2/4 2yy'=-x/4 ?

myininaya (myininaya):

that is actually really really good now solve for y'

myininaya (myininaya):

oops and that is 2yy'=-x/2 right?

OpenStudy (anonymous):

y'=x/4y

OpenStudy (anonymous):

yeah u right its 2 not 4

myininaya (myininaya):

and we have y'=-x/(4y)

myininaya (myininaya):

great

myininaya (myininaya):

now we want to find the tangent at the point \[(\sqrt{2},\frac{1}{\sqrt{2}})\] So you found the formula for the general slope of our ellipse

myininaya (myininaya):

plug this point into find our slope m

OpenStudy (anonymous):

yeah so now if we put x,y values its clear that y' at p=- 4

myininaya (myininaya):

then we could just use point slope form to find the line \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

cool ! i got it ur great ;)

myininaya (myininaya):

wait i'm not getting -4

myininaya (myininaya):

\[y'|_{(\sqrt{2},\frac{1}{\sqrt{2}})}=-\frac{\sqrt{2}}{4 \cdot \frac{1}{\sqrt{2}}}\] is this what you did

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

ohh its -1/2 then ?

myininaya (myininaya):

\[=-\frac{\sqrt{2}}{4} \cdot \frac{1}{\frac{1}{\sqrt{2}}}=-\frac{\sqrt{2}}{4} \cdot \sqrt{2}\] right

myininaya (myininaya):

now you got it

OpenStudy (anonymous):

thanks again !

myininaya (myininaya):

np

myininaya (myininaya):

If you haven't learned implicit differentiation, that will be around the corner. Have fun.

OpenStudy (anonymous):

huh no i learned it , im just refreshing rusty concepts :)

myininaya (myininaya):

ah